OpenStudy (anonymous):

simply x^3/x^2, by first expanding x^2 and x^3

OpenStudy (saifoo.khan):

\[\frac{ x \times x \times x}{x \times x} \to \frac{ \cancel{x} \times \cancel{x} \times x}{\cancel{x} \times \cancel{x}}\]

OpenStudy (saifoo.khan):

Finally you are left with, x.

OpenStudy (anonymous):

\[\frac{x\times x\times x}{x\times x}=\frac{\cancel{x}\cancel{x}x}{\cancel{x}\cancel{x}}=x\]

OpenStudy (anonymous):

i am old and slow!

OpenStudy (saifoo.khan):

LOL.

OpenStudy (across):

or saifoo has a supercomputer for a pc ;P

OpenStudy (saifoo.khan):

LOL